Set up an equation and solve the problem
5 times the square of a number equals 45.
Find the number.
Thanks in advance..

Answer by nerdybill(7384)

5 times the square of a number equals 45.
Find the number.
.
Let x = the number
then
5(x^2) = 45
x^2 = 9
x = sqrt(9)
x = +- 3
that's "plus or minus" three